A unit of length is chosen such that the speed of light in vacuum is unity.what is the distance between sun and the earth in terms of the new unit,if light takes 8 min.and 20 sec .to cover thid distance?

Grade:12

2 Answers

Arun
4 years ago
Given that
The speed of light in vacuum is unity
 
And we want to find the distance between the Sun and the Earth
Use formula
Distance = speed  × time 
Speed is unity = 1 unit /sec
Time = 8 min and 20 sec   = 8 × 60 + 20 sec = 500 sec 
Plug the values we get
 
Distance between Sun and Earth = 1× 500 = 500 units
Khimraj
4 years ago
The speed of light in vacuum © = 1 new unit of length /s
Time taken by light to reach the earth
t = 8 min + 20 s = (8 X 60 + 20) s = 500 s
Distance between the sun and the earth = Speed of light x Time
x = c x t = 1 new unit of length/ s x 500 s = 500 new unit of length